Grade 11Physical Chemistry

1 mole of an organic compound containing C,H and O on complete combustion produces 134.4L of CO2 gas at STP and 108 g h2o. Then molecular formulae is

ujjwal
7 Years agoGrade 11

1 Answer

Ravleen Kaur
6 Years ago

To determine the molecular formula of the organic compound based on the combustion data provided, we need to analyze the products formed during combustion: carbon dioxide (CO2) and water (H2O). Let's break this down step by step.

Step 1: Analyzing the Products

When 1 mole of the organic compound is completely combusted, it produces 134.4 liters of CO2 and 108 grams of H2O. To find the amounts of carbon (C), hydrogen (H), and oxygen (O) in the compound, we'll use the stoichiometry of the products.

Calculating Moles of CO2

First, we can calculate the number of moles of CO2 produced. At standard temperature and pressure (STP), 1 mole of any gas occupies 22.4 liters. Therefore, we can find the moles of CO2 from the volume given:

  • Volume of CO2 = 134.4 L
  • Moles of CO2 = Volume / Molar Volume = 134.4 L / 22.4 L/mol = 6 moles

Finding Carbon Content

Since each mole of CO2 contains one mole of carbon, the moles of carbon produced are equal to the moles of CO2:

  • Moles of C = 6 moles

Step 2: Calculating Moles of H2O

Next, we can determine the number of moles of water produced. The molar mass of water (H2O) is approximately 18 g/mol, so we can calculate the moles of H2O as follows:

  • Moles of H2O = Mass / Molar Mass = 108 g / 18 g/mol = 6 moles

Finding Hydrogen Content

Each mole of water contains two moles of hydrogen, so the total moles of hydrogen are:

  • Moles of H = 6 moles H2O × 2 = 12 moles

Step 3: Determining the Amount of Oxygen

Next, we need to find the moles of oxygen in the original compound. The combustion reaction can be summarized as follows:

  • Complete combustion of CxHyOz produces CO2 and H2O:
  • CxHy + O2 → xCO2 + (y/2)H2O

From our combustion results, we have:

  • 6 moles of CO2 contribute 6 moles of C.
  • 12 moles of H comes from 6 moles of H2O.

Now, in a complete combustion reaction, the total number of moles of oxygen consumed can be calculated from the products:

  • From CO2: 6 moles of CO2 means 6 moles of O (one from each CO2).
  • From H2O: 6 moles of H2O means 3 moles of O (half from each H2O).

Thus, total moles of O used = 6 + 3 = 9 moles.

Calculating Moles of Oxygen in the Compound

The original organic compound had 1 mole; therefore, the number of moles of oxygen in the compound is:

  • Moles of O in the compound = Total O used - O from combustion = 9 - 0 = 0 moles.

Step 4: Establishing the Empirical Formula

Now we have the moles of each element:

  • Carbon: 6
  • Hydrogen: 12
  • Oxygen: 0

The simplest ratio of C:H:O is 6:12:0, which simplifies to 1:2:0. This means the empirical formula is C1H2. However, since the oxygen is absent, we’ll consider the only carbon and hydrogen present.

Step 5: Finding the Molecular Formula

To determine the molecular formula, we need to know the molar mass of the compound, which isn't provided. However, if we assume the empirical formula represents the simplest form, the molecular formula would also be C6H12, as 1 mole of the compound is present.

In summary, the molecular formula of the organic compound combusted is likely C6H12, given the combustion data and the empirical deduction based on the produced gases.